OpenStudy (anonymous):

The sum of the measures of the interior angles of a convex polygon is 2700°. How many sides does the polygon have? A. 15 B. 17 C. 18 D. 19

OpenStudy (mathmale):

Please refer to the following results of my Internet search for "interior angles of a polygon:" https://www.google.com/search?q=interior+angles+of+a+polygon&rlz=1C1CHFX_enUS461US461&oq=interior+angles+of+a+polygon&aqs=chrome..69i57j0l5.4415j0j7&sourceid=chrome&espv=210&es_sm=122&ie=UTF-8

OpenStudy (anonymous):

B... formula: (n-2)180 =(17-2)180 =(15)180 =2700